OpenStudy (anonymous):

simplify and show steps please: (4^(1/3))^2+8/(4^1/3))

OpenStudy (asnaseer):

first simplify each part of this:\[\frac{8}{4^{\frac{1}{3}}}=\frac{2*4}{4^{\frac{1}{3}}}=2*4^{\frac{2}{3}}\]do you understand this step?

OpenStudy (anonymous):

yessir

OpenStudy (asnaseer):

ok, next we can simplify the other part of your equation to get:\[(4^{\frac{1}{3}})^2=4^\frac{2}{3}\]then simply put it all together to get:\[(4^{\frac{1}{3}})^2+\frac{8}{4^{\frac{1}{3}}}=4^\frac{2}{3}+2*4^\frac{2}{3}=3*4^\frac{2}{3}\]

OpenStudy (asnaseer):

do you follow, or do you want more explanation?

OpenStudy (anonymous):

i follow

OpenStudy (asnaseer):

ok - thats good.